OpenStudy (anonymous):

Here's a toughie. At each point (x,y) on the curve the slope equals the square of the distance between the point and the y-axis; the point (-1,2) is on the curve. Find the equation of this function.

OpenStudy (anonymous):

Knowing that the x-value is the perfect measure for the distance from the y-axis, I came up with\[f'(x)=(\left| x \right|)^2\] as the eqn for the slope. But does that make sense...?

OpenStudy (anonymous):

isnt that just: \[y' = x^2 \Rightarrow y = \frac{x^3}{3}\] or am i over simplifying things <.<

OpenStudy (anonymous):

i dont like calculus btw >.<

OpenStudy (anonymous):

Just adjust with with Constant \[2=\frac{1}{3}(-1)^3+C\] \[y=\frac{1}{3}x^3+\frac{7}{3}\]

OpenStudy (anonymous):

Hmm. I guess throwing in those absolute value bars messed me up. Thanks all!

OpenStudy (anonymous):

We don't need absolute value since x^2 is always positive

OpenStudy (anonymous):

y=1x+3